# NTSEBench: Cognitive Reasoning Benchmark for Vision Language Models

### v2

@@ Abstract
Cognitive textual and visual reasoning tasks, including puzzles, series, and analogies, demand the ability to quickly reason, decipher, and evaluate patterns both textually and spatially. Due to extensive training on vast amounts of human-curated data, LLMs and VLMs excel in common-sense reasoning tasks, however still struggle with more complex reasoning that demands deeper cognitive understanding. We introduce NTSEBench, a new dataset designed to evaluate cognitive multi-modal reasoning and problem-solving skills of large models. The dataset contains 2,728 multiple-choice questions, accompanied by a total of 4,642 images, categorized into 26 different types. These questions are drawn from the nationwide NTSE examination in India and feature a mix of visual and textual general aptitude challenges, designed to assess intelligence and critical thinking skills beyond mere rote learning. We establish baselines on the dataset using state-of-the-art LLMs and VLMs. To facilitate a comparison between open source and propriety models, we propose four distinct modeling strategies to handle different modalities—text and images—in the dataset instances.
@@ 1 Introduction
Figure 1: NTSEBench Examples: Shows three samples of textual, direction, and spatial reasoning questions from the proposed dataset. Solutions to these questions are not included here but are provided in the dataset.
@@
The capabilities of LLM models rivals humans on various tasks—question answering(QA), sentiment classification, text generation, visual QA, coding challenges, and mathematical reasoning, to name a few Srivastava et al. (2022); Bubeck et al. (2023). LLMs and Vision Language Models (VLMs) also know as Multimodal Large Language Models(MLLM) have hence become the default benchmark for many zero or few-shot text and vision-based tasks Brown et al. (2020); Wei et al. (2022). Training on vast datasets from diverse domains has enabled LLMs to achieve human-level performance in SAT, GRE, and AP exams, and on platforms like LeetCode Achiam et al. (2023); Touvron et al. (2023); Dubey et al. (2024).
@@
The results from King (2023) indicate that LLMs excel in tasks involving textual reasoning, such as comprehension, analogies, and identifying opposites, but struggle with other types of questions. While Multi-modal Large Language models have demonstrated remarkable performance in many tests of human intelligence, yet they still fall short of human baselines in tasks that require critical and logical thinking, such as commonsense-, numerical- and scientific-reasoning, puzzles, and analogies Most existing visual and multi-modal reasoning datasets are domain-specific, concentrating on fields like science, engineering, and medicine Yue et al. (2024); Zhang et al. (2024b); Sun et al. (2024). These datasets primarily focus on tasks related to concrete scenarios or specific domains, often requiring domain knowledge and rote learning for high performance. However, they do not adequately assess intelligence as a function of cognitive/critical reasoning skills such as spatial recognition, visual puzzle solving, abstract reasoning, or pattern recognition. In this study, we aim to address this research gap by introducing a novel benchmark dataset, NTSEBench, specifically created to evaluate the complex visual, textual, and multi-modal cognitive reasoning capabilities of large deep learning models such as LLMs and VLMs. Given the potential of LLM/VLMs to replace human jobs in the future, our dataset aims to answer the following question: Do LLM/VLMs possess some form of general human intelligence?. Examples of questions from the proposed dataset are shown in Figure 1.
@@
NTSEBench is dedicated to establishing a benchmark for testing capabilities that do not rely on domain-specific knowledge or rote learning. Its primary contribution lies in evaluating the innate problem-solving skills inherent in human cognitive development and isolating where models lack by presenting well-categorized data. It comprises of questions sourced from the nationwide talent search examination (NTSE) conducted in India. These questions can be presented in text format, visual format, or both (multi-modal). We evaluate the performance of recent LLMs and VLMs, including both Proprietary Achiam et al. (2023); Reid et al. (2024) and open source models Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024a); Wang et al. (2023a); Bai et al. (2023); Li et al. (2024); Lu et al. (2024) on our dataset and perform in-depth error analysis of LLM responses to pinpoint areas of weakness and evaluate their overall robustness. Our work makes the following contributions:
@@
  • NTSEBench, a dataset to evaluate complex textual, visual and multi-modal cognitive reasoning capabilities with 2,728 questions in 26 different problem categories.
@@
  • Assess performance using various modeling strategies to effectively handle multimodal (multi image inputs as well) inputs for reasoning-based questions.
@@ 2 NTSEBench Benchmark
NTSE Exam. The National Talent Search Examination (NTSE), administered by the National Council of Educational Research and Training (NCERT) in India since 1963, is a nationwide exam for secondary-grade students. The exam consists of two sections designed to assess a wide range of analytical skills: the Mental Ability Test (MAT) and the Scholastic Aptitude Test (SAT). The MAT section evaluates students’ general aptitude, critical thinking, logical and spatial reasoning, and analytical problem-solving skills (for both textual and visual problems). In contrast, the SAT assesses their domain-specific knowledge in science and mathematics. All questions in the NTSE are multiple-choice (MCQ) with one correct option. Questions and options can be text/image or a combination of both, i.e., multi-modal. We aim to create a dataset focused on cognitive reasoning abilities (MAT-type questions).
Cognitive Reasoning. Cognitive understanding in the context of NTSEBench refers to the ability to process information, recognize patterns, draw inferences, and solve problems using critical, logical and analytical reasoning. This aligns with fundamental concepts in cognitive science, such as problem-solving, pattern recognition, and inferential reasoning Wang and Chiew (2010). It encompasses advanced reasoning skills typically found in a small subset of the population, generally individuals with very high IQs. To emphasize this distinction, we use the term cognitive reasoning for our dataset, differentiating it from common sense reasoning tasks Sakaguchi et al. (2021); Talmor et al. (2019). NTSEBench assesses these reasoning abilities through diverse question categories, each targeting a different cognitive dimension:
  • Pattern Recognition: Categories like Series (Numerical, Alphabetical, Alphanumeric), Missing Character, Non-Verbal Series, and Dot Problem test the ability to identify and extend patterns, which is crucial for understanding sequences and predicting outcomes.
  • Logical Deduction: Blood Relation, Syllogisms, Statement & Conclusions, and Data Sufficiency categories focus on making inferences and drawing conclusions based on the given information, reflecting the core of logical reasoning.
  • Spatial Reasoning: Direction Sense, Cube and Dice, Paper Folding & Cutting, and Embedded Figure assess the ability to visualize and manipulate objects in space, which is essential for understanding spatial relationships.
@@
  • Quantitative Analysis: Number and Ranking, Mathematical Operations, Time and Clock, and Figure Partition test numerical problem-solving skills and the ability to manage quantitative data effectively.
@@
  • Contextual Interpretation: Looking specific details, instructions, or constraints that are critical to understanding and solving the problem.
@@
Unlike other benchmarks that focus on specific academic domains or individual cognitive dimensions as stated above, NTSEBench emphasizes a wide range of cognitive skills, offering a more comprehensive assessment of reasoning abilities. We have expanded on our proposed dataset’s global relevance in Appendix  A.1.
@@ 2.1 Dataset Sources
We created the dataset using past NTSE papers and solutions from Resonance. Additionally, we used NTSE preparation materials, such as a reference book titled A Modern Approach to Verbal and Non-Verbal Reasoning which includes additional logical reasoning problems. We also incorporated content from another book titled Study Guide for NTSE to construct our dataset. The question extraction process is detailed in Appendix A.2.1. The full dataset will be released upon acceptance, with example questions in Figure 1 and Appendix B.
Problem Categories: NTSEBench encompasses several problem categories, each designed to test a distinct set of skills. Questions from these categories frequently appear in NTSE exams year after year. A detailed description of each category is done in Appendix Table 4 and examples shown in Appendix section. More dataset related information is present in Appendix section A.2.
@@ 3 Models: LLMs and VLMs
Modelling Strategies.
Evaluating the reasoning abilities of large language models (LLMs) with text-based questions is straightforward. For vision-language models (VLMs), reasoning with vision-text questions is generally straightforward. Some API access model models, like GPT-4o and Gemini, support multi-image inputs, while many others do not (open-source models like LLaVA-OneVision and Ovis are emerging with this capability). To address these task-specific and input-related dependencies, we propose four strategies to fairly evaluate the reasoning abilities of both open-source and API-based models.
  • Standard QA. For instances where question type($J$) for questions($Q$), options($O$) and solutions($S$) is text($T$), we use a standard text-based QA model like GPT3.5-Turbo or Llama3-70b AI@Meta (2024) or Mixtral8x7b Jiang et al. (2024a).
@@
  • Interleaved model. In this approach, we integrate text with multiple images to create an interwoven context. This method involves placing related textual and visual elements in proximity, enhancing the model’s ability to draw connections between them.
@@
Example inputs for each of the above modelling strategy proposed are shown in Appendix Figure 3.
@@
Implementation Details. We evaluate NTSEBench using multiple open and closed source LLMs Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023) and VLMs Reid et al. (2024); Bai et al. (2023); Dong et al. (2024); Wang et al. (2023a). We used low temperature setting to promote reproducibility. Details on models and hyperparameter can be found in Appendix  A.5.
In the few-shot settings, the questions were sorted by solution length, then an annotator picked the three examples with the most comprehensive responses.
Answer Extraction and Evaluation. The answer extraction module uses a rule-based approach to identify the correct option from the response Gupta et al. (2023). Responses where the answers can’t be extracted by module are evaluated by humans. We used percentage accuracy as our metric for evaluation across all the models. If the model chooses the right option in the MCQ question, label is set of True, else label is False. Option Bias Ablation. We conducted option shuffling experiments to assess whether model performance is influenced by the position of the correct option and to detect potential bias. Four dataset variations were created, placing the correct answer in positions 1, 2, 3, and 4, respectively. Each experiment was run in three rounds, and results were averaged to ensure robustness.
@@ 4 Results and Analysis
1. How well do LLMs and VLMs perform on advanced textual, visual, and multi-modal reasoning questions? How challenging is cognitive reasoning for the current state-of-the-art models?
2. Are proprietary models superior to open-source models, and by what margin? Are there specific categories where proprietary models significantly outperform open-source models?
3. Do different modeling strategies affect a model’s accuracy? Does OCR impact the reasoning accuracy of models?
4. Does shuffling the options in questions impact model performance, indicating a potential option bias in large models?
Results for text only questions using Standard QA and Image Only (implicit OCR) modelling strategy are shown in Table 2. For Standard QA strategy, we report results for zero shot setting in all 10 models listed in the appendix  A.5. For Image Only modelling results are reported for five models which support vision or multi-modal inputs for zero shot results. Columns in the table refers to different types of questions in NTSEBench such as SER (series), ALP (alphabet Test), ODO (odd one out), ANA (analogy), COD (coding-decoding), NUM (number and ranking), BLR (blood relation), MTO (mathematical operations), PUZ (puzzle test), SYL (syllogisms), STC (statement & conclusions) and DAT (data sufficiency).
@@
Table 2: Text Only Question Results. Zero-shot and Few-shot performance of different models across various text-only categories. We report results using two different modelling strategies Image Only and Standard QA. italics font for propriety models, i.e., money or API access is required to run these models. The # is due to the category’s solution contains images thus restricting few shot on text-only models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
@@
Table 3: Multi-modality Question Results: Zero-shot and Few-shot performance of different VLM models across various Text+Vision categories. We report results using all three different modelling strategies proposed to handle multimodality data, i.e., Interleaved, Image Only and Standard VQA. italics font for proprity models, i.e., money or API access is required to run these models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
@@
Results for multi-modality question are reported in Table 3. Results are reported using three different modelling strategies listed in the section above namely Interleaved, Image Only and Standard VQA. Columns in the table refer to different categories of visual text reasoning questions in NTSEBench such as DIR (direction sense), VEN (venn diagrams), TIM (time and clock), MIS (missing character), NVS (non-verbal series), NVO (non-verbal odd one out), NVA (non-verbal analogy), INC (incomplete figure), MIR (mirror, water and images), CUB (cube and dice), PAP (paper folding & cutting), EMB (embedded figure), FIG (figure partition), DOT (dot problem). Option bias ablation results for Gemini-1.5-pro and GPT-4o models are shown in Appendix Tables 8,  9,  10 and  11.
Proprietary models outperform Open Source models. From results in table 2 and 3, we can observe that proprietary models, such as Gemini Pro 1.5, GPT-4o and o1-preview, outperform other open source models in nearly every category question category. Especially o1-preview, which operates by utilizing internal chain-of-thought reasoning, allowing it to analyze complex problems step-by-step before arriving at a solution. Proprietary models demonstrate nearly double the accuracy of open-source models on NTSEBench questions, both in text-based and multi-modal tasks, across all modeling strategies outlined in the previous section. Notably, Gemini Pro 1.5 consistently outperforms GPT-4o across most strategies, excelling in both text and multi-modal question types. For open-source models, LLaVA-OneVision-72b-ov-chat excels in text and multi-modality tasks, achieving SOTA results in some vision categories. Ovis1.6-Gemma2-9B, while smaller, outperforms GPT-4o and LLaVA-OneVision in several vision categories, though its overall performance is still lower.
@@
Modelling Strategy is important. For text-only questions, the Standard QA strategy clearly outperforms Image Only modeling. Introducing burden of doing OCR on top of reasoning tends to confuse models or exacerbate the difficulty of the task. This effect is particularly noticeable with smaller open-source models, which struggle to accurately extract characters and integrate them into context. However, proprietary models like GPT-4o and Gemini-Pro still achieve superior results using Image Only compared to open-source models employing Standard QA or text-only processing alone. Few-shot results present a mixed picture: while models like Mixtral and GPT-4o show improved performance with added exemplars, others experience a significant decline. Tables 2 and 3 show that proprietary models generally experience a smaller performance drop than open-source models in such scenarios.
@@
Interleaving text and images performs better than Standard VQA and Image Only strategy for most categories. This underscores the importance of presenting text and images separately and in a more detailed manner, providing appropriate context, and treating the image and text as distinct entities. Our results in Table 3 show that this approach significantly improves Vision-and-Language Models (VLMs), as demonstrated by the superior performance of open-source models like LLaVA-OneVision-72b and Ovis-9b, which handle interleaved text and images more effectively than others. Table 3 shows that few-shot prompting consistently underperforms compared to zero-shot for all VLM models, suggesting that adding exemplars may confuse large models, hindering their focus on logical reasoning.
@@
Multi-modal reasoning is significantly harder. Comparing the best and worst performing models in Tables 2 and 3, it’s evident that multi-modal reasoning is considerably more challenging than textual reasoning for current state-of-the-art VLMs. Multi-modal questions see less than 45% accuracy, while the best model for textual reasoning, o1-preview, exceeds 80%. Even fast models like Gemini 1.5 pro achieve over 60% accuracy in textual reasoning, highlighting the gap between VLMs and LLMs and the need for better architectures and datasets for VLMs.
@@
Question category analysis. The results from Table 2 reveal that while LLMs generally perform better on the text-only subset of NTSEBench, the high standard deviation (11.04) indicates significant variability in model performance across different question types. This variability may stem from some overlap between NTSEBench and other open-source datasets, suggesting that there are still areas where LLMs exhibit limitations in reasoning capabilities. This is especially evident in Alphabet Test(ALP) category and Mathematical Operations(MTO) category, where the accuracy is more than two standard deviations away from the mean accuracy of the model. This could also be attributed to the potential difficulty of these question types however that analysis has been left for future work.
@@
VLMs have shown poor performance across all categories for multi-modal questions, with the best-performing model achieving correct answers only 42% of the time. Even the standard deviation for accuracy of the best model is 9, indicating that VLMs struggle more with certain question categories than others. Specifically, we observe that VLMs perform notably poorly on categories such as DOT (Dot Problems), NVS (Non-Verbal Series), and NVO (Non-Verbal Odd One Out). These categories require identifying correlations or patterns between multiple images or recognizing emerging patterns in a sequence of images. This task is akin to identifying similar and evolving patterns in different parts of an image and predicting the next possible pattern. While vision models excel at recognizing existing patterns, they struggle with predicting new patterns.

### v3

@@ Abstract
Cognitive textual and visual reasoning tasks, including puzzles, series, and analogies, demand the ability to quickly reason, decipher, and evaluate patterns both textually and spatially. Due to extensive training on vast amounts of human-curated data, large language models (LLMs) and vision language models (VLMs) excel in common-sense reasoning tasks, but still struggle with more complex reasoning that demands deeper cognitive understanding. We introduce NTSEBench, a new dataset designed to evaluate cognitive multimodal reasoning and problem-solving skills of large models. The dataset contains 2,728 multiple-choice questions, accompanied by a total of 4,642 images, spanning 26 categories. These questions are drawn from the nationwide NTSE examination in India and feature a mix of visual and textual general aptitude challenges, designed to assess intelligence and critical thinking skills beyond mere rote learning. We establish baselines on the dataset using state-of-the-art LLMs and VLMs. To facilitate a comparison between open-source and propriety models, we propose four distinct modeling strategies to handle different modalities—text and images—in the dataset instances.
@@ 1 Introduction
Figure 1: NTSEBench Examples: Three samples of textual, direction, and spatial reasoning questions from the proposed dataset. Solutions to these questions are not included here but are provided in the dataset.
@@
The capabilities of LLM models rivals humans on various tasks—question answering (QA), sentiment classification, text generation, visual QA, coding challenges, and mathematical reasoning, to name a few Srivastava et al. (2022); Bubeck et al. (2023). LLMs and VLMs (also known as Multimodal Large Language Models, MLLMs) have hence become the default benchmark for many zero or few-shot text and vision-based tasks Brown et al. (2020); Wei et al. (2022). Training on vast datasets from diverse domains has enabled LLMs to achieve human-level performance in SAT, GRE, and AP exams, and on platforms such as LeetCode Achiam et al. (2023); Touvron et al. (2023); Dubey et al. (2024).
@@
The results from King (2023) indicate that LLMs excel in tasks involving textual reasoning such as comprehension, analogies, and identifying opposites but struggle with other types of questions. Multimodal Large Language models have demonstrated remarkable performance in many tests of human intelligence, but they still fall short of human baselines in tasks that require critical and logical thinking, such as commonsense-, numerical- and scientific-reasoning, puzzles, and analogies. Most existing visual and multimodal reasoning datasets are domain-specific, concentrating on fields such as science, engineering, and medicine Yue et al. (2024); Zhang et al. (2024b); Sun et al. (2024). These datasets primarily focus on tasks related to concrete scenarios or specific domains, often requiring domain knowledge and rote learning for high performance. However, they do not adequately assess intelligence as a function of cognitive/critical reasoning skills such as spatial recognition, visual puzzle solving, abstract reasoning, or pattern recognition. In this study, we aim to address this research gap by introducing a novel benchmark dataset, NTSEBench, specifically created to evaluate the complex visual, textual, and multimodal cognitive reasoning capabilities of large deep learning models. Examples of questions from the proposed dataset are shown in Figure 1.
@@
NTSEBench is dedicated to establishing a benchmark for testing capabilities that do not rely on domain-specific knowledge or rote learning. Its primary contribution lies in evaluating the innate problem-solving skills inherent in human cognitive development and isolating where models are lacking by presenting well-categorized data. It comprises questions sourced from the Nationwide Talent Search Examination (NTSE) conducted in India. These questions can be presented in text format, visual format, or both (multimodal). We evaluate the performance of recent LLMs and VLMs, including both proprietary Achiam et al. (2023); Reid et al. (2024) and open-source models Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024a); Wang et al. (2023a); Bai et al. (2023); Li et al. (2024); Lu et al. (2024) on our dataset and perform in-depth error analysis of LLM responses to pinpoint areas of weakness and evaluate their overall robustness. Our work makes the following contributions:
@@
  • NTSEBench, a dataset to evaluate complex textual, visual, and multimodal cognitive reasoning capabilities with 2,728 questions in 26 problem categories.
@@
  • Assess performance using various modeling strategies to effectively handle multimodal (multi-image inputs as well) inputs for reasoning-based questions.
Code and dataset for the experiments with NTSEBench are available at https://ntsebench.github.io/.
@@ 2 NTSEBench Benchmark
NTSE Exam. The National Talent Search Examination (NTSE), administered by the National Council of Educational Research and Training (NCERT) in India since 1963, is a nationwide exam for secondary-grade students. The exam consists of two sections designed to assess a wide range of analytical skills: the Mental Ability Test (MAT) and the Scholastic Aptitude Test (SAT). The MAT section evaluates students’ general aptitude, critical thinking, logical and spatial reasoning, and analytical problem-solving skills (for both textual and visual problems). In contrast, the SAT assesses their domain-specific knowledge in science and mathematics. All questions in the NTSE are multiple-choice (MCQ) with one correct option. Questions and options can be text or image or a combination of both, i.e., multimodal. Our aim is to create a dataset focused to test cognitive reasoning abilities (or MAT-type questions).
Cognitive Reasoning. Cognitive understanding in the context of NTSEBench refers to the ability to process information, recognize patterns, draw inferences, and solve problems using critical, logical, and analytical reasoning. This aligns with fundamental concepts in cognitive science, such as problem-solving, pattern recognition, and inferential reasoning Wang and Chiew (2010). It encompasses advanced reasoning skills typically found in a small subset of the population, generally individuals with very high IQs. To emphasize this distinction, we use the term cognitive reasoning for our dataset, differentiating it from common sense reasoning tasks Sakaguchi et al. (2021); Talmor et al. (2019). NTSEBench assesses these reasoning abilities through diverse question categories, targeting a different cognitive dimension:
  • Pattern Recognition: Categories such as Series (Numerical, Alphabetical, Alphanumeric), Missing Character, Non-verbal Series, and Dot Problem test the ability to identify and extend patterns, which is crucial for understanding sequences and predicting outcomes.
  • Logical Deduction: Blood Relation, Syllogisms, Statement and Conclusions, and Data Sufficiency categories focus on making inferences and drawing conclusions based on the given information, reflecting the core of logical reasoning.
  • Spatial Reasoning: Direction Sense, Cube and Dice, Paper Folding and Cutting, and Embedded Figure assess the ability to visualize and manipulate objects in space, which is essential for understanding spatial relationships.
@@
  • Quantitative Analysis: Number and Ranking, Mathematical Operations, Time and Clock, and Figure Partition test numerical problem-solving skills and the ability to manage quantitative data.
@@
  • Contextual Interpretation: Looking for specific details, instructions, or constraints that are critical to understanding and solving the problem.
@@
Unlike other benchmarks that focus on specific academic domains or individual cognitive dimensions as stated above, NTSEBench emphasizes a wide range of cognitive skills, offering a more comprehensive assessment of reasoning abilities.
@@ 2.1 Dataset Sources
We created the dataset using past NTSE papers and solutions from Resonance. Additionally, we used NTSE preparation materials, such as a reference book titled A Modern Approach to Verbal and Non-Verbal Reasoning, which includes additional logical reasoning problems. We also incorporated content from another book titled Study Guide for NTSE to construct our dataset. The question extraction process is detailed in Appendix A.1.1. The example questions of the released test dataset in Figure 1 and Appendix B.
Problem Categories: NTSEBench encompasses several problem categories, each designed to test a distinct set of skills. Questions from these categories frequently appear in NTSE exams year after year. A detailed description of each category is done in Appendix Table 5 and examples are shown in Appendix section  B. More dataset related information is present in Appendix section A.1.
@@
Table 1 shows a skewed distribution across various question categories. Notably, textual categories such as the Alphabet Test (ALP) and Mathematical Operations (MTO) contain 94 and 99 examples, respectively. In contrast, many vision-based categories are more challenging and typically include between 80 and 100 examples. For instance, the Non-Verbal Analogy (NVA) category, one of the most difficult, comprises 100 examples. Although this skewed distribution could impact model performance, exploring its effects is beyond the scope of this manuscript and is left for future work.
Modality Variations.
Since NTSEBench has multimodal questions, options, and solutions, we have results in eight combinations of modality types that can occur for question-options-solution triplet. Table 2 shows the count of each triplet option. NTSEBench has 1199 textual questions and the remaining 1529 are multimodal questions.
Question Options Solutions # Samples
1199
381
70
18
330
126
403
201
2.2 The Global Relevance of the NTSE Exam for AI
The NTSE exam, despite being conducted in India, holds significant relevance for the global AI community due to its unique focus on cognitive reasoning abilities rather than domain-specific knowledge. The NTSE’s diverse question categories as described in Table  1( and Appendix table  5), assess various cognitive dimensions, offering a robust framework for evaluating AI models’ capacity to process information, recognize patterns, and solve problems across different domains. This emphasis on cognitive dimensions aligns with the pursuit of Artificial General Intelligence (AGI), making the NTSE exam’s insights and challenges applicable to AI research and development on a global scale.
@@ 3 Models: LLMs and VLMs
Modeling Strategies.
Evaluating the reasoning abilities of large language models (LLMs) with text-based questions is straightforward. For vision-language models (VLMs), reasoning with vision-text questions is generally not straightforward. Some API access model models, such as GPT-4o Achiam et al. (2023) and Gemini Reid et al. (2024), support multi-image inputs, but many others do not (open-source models like LLaVA-OneVision Li et al. (2024) and Ovis Lu et al. (2024) are emerging with this capability). To address these task-specific and input-related dependencies, we propose four strategies to fairly evaluate the reasoning abilities of both open-source and proprietary models.
  • Standard QA. For instances where question type($J$) for questions($Q$), options($O$) and solutions($S$) is text($T$), we use a standard text-based QA model such as GPT3.5-Turbo or Llama3-70b AI@Meta (2024) or Mixtral8x7b Jiang et al. (2024a).
@@
  • Interleaved model. In this approach, we integrate text with multiple images to create an interwoven context. This method involves placing related textual and visual elements in proximity, enhancing the model’s ability to draw connections.
@@
Example inputs for each of the above modeling strategy proposed are shown in Appendix Figure 3.
@@
Implementation Details. We evaluate NTSEBench using multiple open-source and proprietary LLMs Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023) and VLMs Reid et al. (2024); Bai et al. (2023); Dong et al. (2024); Wang et al. (2023a). We used a low temperature setting to promote reproducibility. Details on models used and their hyperparameters can be found in Appendix  A.4.
In the few-shot settings, the questions were sorted by solution length, and then an annotator picked the three examples with the most comprehensive responses.
Answer Extraction and Evaluation. The answer extraction module uses a rule-based approach to identify the correct option from the response Gupta et al. (2023). Responses where the answers cannot be extracted by module are evaluated by humans. We used percentage accuracy as our metric for evaluation across all the models. If the model chooses the right option in the MCQ question, label is set as True, else label is False.
Option Bias Ablation. We conducted option shuffling experiments to assess whether model performance is influenced by the position of the correct option and to detect potential bias. Four dataset variations were created, placing the correct answer in positions 1, 2, 3, and 4, respectively. Each experiment was run in three rounds, and results were averaged to ensure robustness.
@@ 4 Results and Analysis
  • How well do LLMs and VLMs perform on advanced textual, visual, and multimodal reasoning questions? How challenging is cognitive reasoning for the current state-of-the-art models?
  • Are proprietary models superior to open-source models, and if so then by what margin? Are there specific categories where proprietary models significantly outperform open-source models?
  • Do different modeling strategies affect a model’s accuracy? Does OCR impact the reasoning accuracy of models?
  • Does shuffling the options in questions impact model performance, indicating a potential option bias in large models?
Results for text-only questions using Standard QA and Image Only (implicit OCR) modeling strategy are shown in Table 3. For the Standard QA strategy, we report results for zero shot setting in all 10 models listed in the appendix  A.4. For Image Only modeling, results are reported for five models that support vision or multimodal inputs for zero shot results. Columns in the table refers to different types of questions in NTSEBench such as SER (series), ALP (alphabet Test), ODO (odd one out), ANA (analogy), COD (coding-decoding), NUM (number and ranking), BLR (blood relation), MTO (mathematical operations), PUZ (puzzle test), SYL (syllogisms), STC (statement and conclusions), and DAT (data sufficiency).
@@
Table 3: Text Only Question Results. Zero-shot and few-shot performance of different models across various text-only categories. We report results using two modeling strategies Image Only and Standard QA. italics font for propriety models, i.e., money or API access is required to run these models. The # is due to the category’s solution contains images thus restricting few shot on text-only models. Note: (*) In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent, e.g., thinking it as CAPTCHA.
@@
Table 4: Multi-modality Question Results: Zero-shot and few-shot performance of different VLMs across various Text+Vision categories. We report results using all three different modelling strategies proposed to handle multimodality data, i.e., Interleaved, Image Only and Standard VQA. italics font for proprietary models, i.e., money or API access is required to run these models. Note: (*) In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent, e.g., thinking it as CAPTCHA.
@@
Results for the multimodality questions are reported in Table 4. Results are reported using three modeling strategies listed in the section above namely Interleaved, Image Only and Standard VQA. Columns in the table refer to different categories of visual text reasoning questions in NTSEBench such as DIR (direction sense), VEN (venn diagrams), TIM (time and clock), MIS (missing character), NVS (non-verbal series), NVO (non-verbal odd one out), NVA (non-verbal analogy), INC (incomplete figure), MIR (mirror, water and images), CUB (cube and dice), PAP (paper folding and cutting), EMB (embedded figure), FIG (figure partition), DOT (dot problem). Option bias ablation results for Gemini-1.5-pro and GPT-4o models are shown in Appendix Tables 8,  9,  10, and  11. Proprietary models outperform open-source models. From results in Table 3 and 4, we can observe that proprietary models, such as Gemini Pro 1.5, GPT-4o and o1-preview, outperform other open-source models in nearly every question category, especially o1-preview, which operates by utilizing internal chain-of-thought reasoning, allowing it to analyze complex problems step-by-step before arriving at a solution. Proprietary models demonstrate nearly double the accuracy of open-source models on NTSEBench questions, both in text-based and multimodal tasks, across all modeling strategies outlined in the previous section. Notably, Gemini Pro 1.5 consistently outperforms GPT-4o across most strategies, excelling in both text and multimodal question types.
For open-source models, LLaVA-OneVision-72b-ov-chat excels in text and multimodality tasks, achieving SOTA results on tasks such as Non-verbal Odd One Out and Embedded Figure, also matching or surpassing proprietary models in several other categories. Ovis1.6-Gemma2-9B, although smaller, outperforms GPT-4o and LLaVA-OneVision in several vision categories, but its overall performance is still lower.
@@
Modeling Strategy is important. For text-only questions, the Standard QA strategy clearly outperforms Image Only modeling. Introducing the burden of doing OCR on top of reasoning tends to confuse models or exacerbate the difficulty of the task. This effect is particularly noticeable with smaller open-source models, which struggle to accurately extract characters and integrate them into context. However, proprietary models such as GPT-4o and Gemini-Pro still achieve superior results using Image Only compared to open-source models employing Standard QA or text-only processing alone. Few-shot results present a mixed picture: where as models like Mixtral and GPT-4o show improved performance with added exemplars, others experience a significant decline. Tables 3 and 4 show that proprietary models generally experience a smaller performance drop than open-source models in such scenarios.
@@
Interleaving text and images performs better than Standard VQA and Image Only strategy for most categories. This underscores the importance of presenting text and images separately and in a more detailed manner, providing appropriate context, and treating the image and text as distinct entities. Our results in Table 4 show that this approach significantly improves VLM results, as demonstrated by the superior performance of open-source models such as LLaVA-OneVision-72b and Ovis-9b, which handle interleaved text and images more effectively than others. Table 4 also shows that few-shot prompting consistently underperforms compared to zero-shot for all VLM models, suggesting that adding exemplars may confuse VLMs, hindering their focus on logical reasoning.
@@
Multimodal reasoning is significantly harder. Comparing the best and worst performing models in Tables 3 and 4, it is evident that multimodal reasoning is considerably more challenging than textual reasoning for current state-of-the-art models. Multimodal questions see less than 45% accuracy, whereas the best model for textual reasoning, o1-preview, exceeds 80%. Even fast models such as Gemini 1.5 pro achieve over 60% accuracy in textual reasoning, highlighting the gap between VLMs and LLMs and the need for better architectures and datasets for VLMs.
@@
Question category analysis. The results from Table 3 reveal that even though LLMs generally perform better on the text-only subset of NTSEBench, the high standard deviation (11.04) indicates significant variability in model performance across different question types. This variability may stem from some overlap between NTSEBench and other open-source datasets, suggesting that there are still areas where LLMs exhibit limitations in reasoning capabilities. This is especially evident in Alphabet Test (ALP) category and Mathematical Operations (MTO) category, where the accuracy is more than two standard deviations away from the mean accuracy of the model. This could also be attributed to the potential difficulty of these question types; however, that analysis has been left for future work.
@@
VLMs have shown poor performance across all categories for multimodal questions, with the best-performing model achieving correct answers only 42% of the time. Even the standard deviation for accuracy of the best model is 9, indicating that VLMs struggle more with certain question categories than others. Specifically, we observe that VLMs perform notably poorly on categories such as DOT (Dot Problems), NVS (Non-verbal Series), and NVO (Non-verbal Odd One Out). These categories require identifying correlations or patterns between multiple images or recognizing emerging patterns in a sequence of images. This task is akin to identifying similar and evolving patterns in different parts of an image and predicting the next possible pattern. Although vision models excel at recognizing existing patterns, they struggle with predicting new patterns.